# leftmargin=1em, rightmargin=1em, indentfirst=false

### v2

@@ 4.2 Rule-like propositional formula transformation
To achieve this, we first transform $\psi$ by a formula set $\Gamma_{\psi}^{\prime}=\{D_{i}\Rightarrow C_{j}\mid i\in[m],j\in[n]\}$. Then, for each single formula $\psi_{i,j}=D_{i}\Rightarrow C_{j}=l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}\Rightarrow l_{j1}^{C}\vee l^{C}_{j2}\vee\cdots\vee l^{C}_{jn_{j}}$, we reformulate it into the following set of implicant formula:
@@
Consider a DeNF $\psi=(a_{1}\wedge b_{1})\vee a_{2}\Rightarrow(c_{1}\vee d_{1})\wedge c_{2}$. We first transform $\psi$ into an equivalent formula set $\Gamma_{\psi}^{\prime}=\{\psi_{1,1},\psi_{1,2},\psi_{2,1},\psi_{2,2}\}$, where $\psi_{1,1}=a_{1}\wedge b_{1}\Rightarrow c_{1}\vee d_{1}$, $\psi_{1,2}=a_{1}\wedge b_{1}\Rightarrow c_{2}$, $\psi_{2,1}=a_{2}\Rightarrow c_{1}\vee d_{1}$, and $\psi_{2,2}=a_{2}\Rightarrow c_{2}$. For each single formula in $\Gamma_{\psi}^{\prime}$, we then reformulate it as follows: $\Gamma_{\psi_{1,1}}^{\prime}=\{a_{1}\wedge b_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge b_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{1,2}}^{\prime}=\{a_{1}\wedge b_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,1}}^{\prime}=\{a_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,2}}^{\prime}=\{a_{2}\wedge c_{2}\}$, and finally obtain the set of rule-like propsitional formulas as $\Gamma_{\psi}=\Gamma_{\psi_{1,1}}^{\prime}\cup\Gamma_{\psi_{1,2}}^{\prime}\cup\Gamma_{\psi_{2,1}}^{\prime}\cup\Gamma_{\psi_{2,2}}^{\prime}$.
@@ 4.3 Forward chaining
We next illustrate the graph construction and the corresponding FC procedure for the rule set $\Gamma_{\psi}=\{\psi_{1},\psi_{2}\}$ under the Level-2 interpretation given in Example 1. We first obtain all the literals as $\texttt{Lit}_{\mathcal{R}}=\{p_{0},p_{1},p_{2},p_{3},p_{4}\}$ with $\texttt{Lit}^{\uparrow}=\{p_{0},p_{2}\}$ and $\texttt{Lit}^{\downarrow}=\{p_{1}\}$, construct the literal nodes correspondingly, and construct the LHS node set as $\{p_{0}\&p_{2},p_{1}\&p_{2}\}$. Then, directed edges are added as follows: $p_{0}\rightarrow p_{0}\&p_{2}$, $p_{2}\rightarrow p_{0}\&p_{2}$, $p_{0}\&p_{2}\rightarrow p_{3}$, $p_{1}\rightarrow p_{1}\&p_{2}$, $p_{2}\rightarrow p_{1}\&p_{2}$, $p_{1}\&p_{2}\rightarrow p_{4}$, resulting an intial graph. Next, we check the logic consistency. For the LHS node $p_{0}\&p_{2}$, since $\{p_{0},p_{2}\}\subseteq\texttt{Lit}^{\uparrow}$, we can update $\texttt{Lit}^{\uparrow}$ as $\texttt{Lit}^{\uparrow}=\texttt{Lit}^{\uparrow}\cup p_{3}=\{p_{0},p_{2},p_{3}\}$ and $\texttt{Lit}^{\downarrow}=\texttt{Lit}^{\downarrow}\cup\neg p_{3}=\{p_{1},\neg p_{3}\}$, where “$p_{3}={\sf True}$” is the new knowledge derived. For the LHS node $p_{1}\&p_{2}$, since $p_{1}\notin\texttt{Lit}^{\uparrow}$, there is no update happens to $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\uparrow}$; consequently, no new knowledge is inferred.
@@ 5 Experiments
To evaluate the effectiveness of RvLLM, we apply it to the runtime verification of LLMs across three representative tasks that require domain-specific knowledge: violation detection against Singapore Rapid Transit Systems Act [1], numerical comparison, and inequality solving. For a comprehensive evaluation, we include both SOTA and non-SOTA LLMs as our benchmark, including Qwen 2.5 (max, plus, turbo, 7B, 14B, 32B, 72B), GPT (4.1, 4.1 mini, 4.1 nano), Gemini 2.0 (Flash, Flash Lite), and DeepSeek-V3. All experiments are conducted on a machine with an Intel (R) Xeon(R) w7-2475X processor. A dedicated guideline for designing interpretation abstraction prompts for the perception agent is provided in the appendices, along with additional experimental details and descriptions of the datasets and specifications used throughout the experiments.
@@ 5.1 Case study 1: violation detection against Singapore Rapid Transit Systems Act
For this case study, we apply RvLLM under the Level-1 interpretation to evaluate the LLM’s responses in detecting violations within contextual scenarios against the Singapore Rapid Transit Systems Act. In this capacity, RvLLM can be regarded as a complementary mechanism to enhance the LLM’s ability in detecting the law violations in this study. We conduct experiments using contextual scenarios derived from [53], consisting of 304 cases in total–280 labeled as involving violations (unsafe) and 24 as not (safe). To incorporate relevant domain knowledge, we carefully encode 31 of the 53 regulations from Singapore Rapid Transit System Act into ESL specifications. It took a research scientist three days to interpret the law and develop corresponding specifications–a one-time effort applicable for verifying any LLM application against these regulations. In this study, the same model serves both as the target LLM and as the perception agent.
@@
Table 1: Violation detection results of LLMs with and without RvLLM against the Singapore Rapid Transit Systems Act [1].
@@
Qwen max [56] 56.2% 91.3% 86.1% 91.3% 29.9% $\uparrow$ 0 3.00
Qwen plus [56] 58.4% 1 81.9% 1 23.5% $\uparrow$ 0 5.32
Qwen turbo [56] 39.3% 1 75% 1 35.7% $\uparrow$ 0 1.64
Qwen 2.5 (7B) [56] 16.0% 1 61.2% 87.0% 45.2% $\uparrow$ 13% $\downarrow$ 2.28
Qwen 2.5 (14B) [56] 38.8% 95.7% 80.4% 87.0% 41.6% $\uparrow$ 8.7% $\downarrow$ 3.11
Qwen 2.5 (32B) [56] 56.6% 95.7% 87.5% 95.7% 30.9% $\uparrow$ 0 2.55
Qwen 2.5 (72B) [56] 57.3% 1 80.4% 95.7% 23.1% $\uparrow$ 4.3% $\downarrow$ 2.57
GPT 4.1 [3] 57.7% 95.7% 81.1% 91.3% 23.4% $\uparrow$ 4.3% $\downarrow$ 3.11
GPT 4.1 mini [3] 39.1% 95.7% 65.1% 95.7% 26% $\uparrow$ 0 3.94
GPT 4.1 nano [3] 11.8% 1 30.0% 1 18.2% $\uparrow$ 0 1.82
Gemini 2.0 Flash [21] 37.0% 1 87.2% 82.6% 50.2% $\uparrow$ 17.4% $\downarrow$ 2.69
Gemini 2.0 Flash Lite [21] 55.1% 95.7% 79.7% 95.7% 24.6% $\uparrow$ 0 1.72
DeepSeek-V3 [39] 78.3% 87.0% 94.0% 82.6% 15.7% $\uparrow$ 4.4% $\downarrow$ 15.84
@@ 5.2 Case study 2: numerical comparison problems
Table 2: Runtime verification results of various LLMs on numerical comparison tasks.
@@
Qwen max [56] 93 7 0 5.03
Qwen plus [56] 99 1 0 7.01
Qwen turbo [56] 94 6(2) 0 2.93
Qwen 2.5 (7B) [56] 0 0 100 N/A
Qwen 2.5 (14B) [56] 90 10(7) 0 5.09
Qwen 2.5 (32B) [56] 98 2 0 4.98
Qwen 2.5 (72B) [56] 96 4 0 6.31
GPT 4.1 [3] 98 2 0 4.84
GPT 4.1 mini [3] 92(1) 8 0 5.37
GPT 4.1 nano [3] 90(3) 10 0 3.73
Gemini 2.0 Flash [21] 93 7 0 4.94
Gemini 2.0 Flash Lite [21] 52 44 4 4.35
DeepSeek-V3 [39] 69 31(5) 0 30.87
Results. Table 2 gives the verification results by RvLLM on numerical comparison tasks across various LLMs. Columns Con. and Incon. report the number of cases where RvLLM detects no inconsistencies and where at least one inconsistency is detected, following the process in Figure 2. The values in parentheses in Column Con. represent cases where the target LLM produces an erroneous comparison result on the original comparison question, yet RvLLM fails to identify any inconsistencies–typically due to incomplete interpretation by the perception agents. In contrast, the values in parentheses in Column Incon. corresponds to cases where the target LLM returns a correct comparison result on the original question, but RvLLM indeed detects some inconsistencies. This usually arises from two main reasons: i) incomplete interpretation, and ii) the target LLM producing an incorrect diagnosis on newly-inferred knowledge. Overall, RvLLM effectively identifies erroneous diagnoses by the target LLM in most cases, with exceptions observed in only 3 cases for GPT 4.1 nano and mini. Among all tested models, Gemini 2.0 Flash Lite and DeepSeek-V3 show the weakest diagnostic performance, and Qwen 2.5 (7B) has the worst language understanding capabilities, failing all interpretation abstraction. Given its strong performance in the previous violation detection task, we conjecture that DeepSeek-V3 is better suited for language understanding tasks rather than those requiring rigorous logical reasoning.
@@ 5.3 Case study 3: inequality solving problems
For this case study, we employ RvLLM under Level-1 interpretation to verify the step-by-step reasoning process of LLMs in inequality solving tasks. We compile a dataset of 40 inequality questions sourced from A-Level H2 Mathematics examination papers [2] and carefully design three specifications to serve as domain knowledge that RvLLM uses for verification. These specifications consist of basic algebraic inequality properties taught at the junior college level, targeting three common LLM reasoning errors in inequality solving: incorrect factorization, flawed interval analysis, and omission of endpoint or critical point checks. Given the higher interpretation complexity in this case study and the strong language processing capabilities of DeepSeek-V3, we employ it as the perception agent for all LLMs evaluated and Qwen 2.5 (32B) as an alternative for comparison.
Results. Given the strong performance of RvLLM in detecting true positives, we focus the evaluation on the questions that the target LLMs answered incorrectly. The results are summarized in Table 3. Column 2 lists the total number of questions for which the target LLM produces incorrect solutions. Columns labeled DS-V3 and QW-32B (spanning Columns 3 to 11) show the number of cases where RvLLM successfully identifies inconsistencies based on the corresponding specifications using DeepSeek-V3 and Qwen 2.5 (32B) as the perception agent, respectively. Values in parentheses indicate false positives–cases where the detected inconsistency does not stem from an actual reasoning error but from inaccuracies of the perception agent during the interpretation abstraction process. Column GT gives the ground truth number of incorrect responses attributable to the rule specifications. It is important to note that we define only three specifications as domain knowledge for inequality-solving tasks, which is insufficient to comprehensively cover all reasoning patterns required for such problems. As a result, it is expected that RvLLM fails to detect a portion of the errors due to the incompleteness of the specifications. The final two columns give the true positive rate of RvLLM with three specifications. We find that RvLLM achieves a TPR of up to 52.2% when using DeepSeek-V3 as the perception agent. This result highlights not only the potential of RvLLM in detecting reasoning errors (even with incomplete domain knowledge), but also underscores the importance of employing a powerful perception agent to enhance the overall performance. Note that, although the TPR remains below 50% for most of the models, it can reach significantly higher values when considering only the ground truth cases listed in Column GT, indicating the potential for performance improvement if more domain knowledge with respect to the inequality solving is incorporated.
@@
Qwen max [56] 12 1 1 1 1 0 1 2 2 3 33.3% 25%
Qwen plus [56] 10 1 1 2 0 0 0 0 1 2 10% 20%
Qwen turbo [56] 22 0 0 1 1 0 1 3 2 4 18.2% 9.1%
Qwen 2.5 (7B) [56] 23 2 1 3 7 2 7 3 2 5 52.2% 21.7%
Qwen 2.5 (14B) [56] 18 0 0 1 3 2(1) 3 3(1) 2 3 27.8% 16.7%
Qwen 2.5 (32B) [56] 16 1 0 1 3 2 4 4 2 5 50% 25%
Qwen 2.5 (72B) [56] 15 1 0 1 4 3 4 0 0 1 33.3% 20
GPT 4.1 [3] 5 1 1(1) 1 0 0 0 0 0 0 20% 0
GPT 4.1 mini [3] 2 0 0 0 0 0 0 1 0 1 50% 0
GPT 4.1 nano [3] 10 0 0 0 1 0 1 1 0 1 20% 0
Gemini 2.0 Flash [21] 5 0 0 0 1 0 1 0 0 0 20% 0
Gemini 2.0 Flash Lite [21] 6 1 1 2 0 0 0 0 0 1 16.7% 16.7%

### v3

@@ 3.2 Rule-like propositional formula transformation
To achieve this, we first transform $\psi$ to a formula set $\Gamma_{\psi}^{\prime}=\{D_{i}\Rightarrow C_{j}\mid i\in[m],j\in[n]\}$. Then, for each single formula $\psi_{i,j}=D_{i}\Rightarrow C_{j}=l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}\Rightarrow l_{j1}^{C}\vee l^{C}_{j2}\vee\cdots\vee l^{C}_{jn_{j}}$, we reformulate it as the implicant formula set $\Gamma_{\psi_{i,j}}^{\prime}$, following the idea of unit propagation [zhang1996cient]:
@@
Consider a DeNF $\psi=(a_{1}\wedge b_{1})\vee a_{2}\Rightarrow(c_{1}\vee d_{1})\wedge c_{2}$. We first transform $\psi$ into an equivalent formula set $\Gamma_{\psi}^{\prime}=\{\psi_{1,1},\psi_{1,2},\psi_{2,1},\psi_{2,2}\}$, where $\psi_{1,1}=a_{1}\wedge b_{1}\Rightarrow c_{1}\vee d_{1}$, $\psi_{1,2}=a_{1}\wedge b_{1}\Rightarrow c_{2}$, $\psi_{2,1}=a_{2}\Rightarrow c_{1}\vee d_{1}$, and $\psi_{2,2}=a_{2}\Rightarrow c_{2}$. For each single formula in $\Gamma_{\psi}^{\prime}$, we then reformulate it as follows: $\Gamma_{\psi_{1,1}}^{\prime}=\{a_{1}\wedge b_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge b_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{1,2}}^{\prime}=\{a_{1}\wedge b_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,1}}^{\prime}=\{a_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,2}}^{\prime}=\{a_{2}\wedge c_{2}\}$, and finally obtain the set of rule-like propositional formulae as $\Gamma_{\psi}=\Gamma_{\psi_{1,1}}^{\prime}\cup\Gamma_{\psi_{1,2}}^{\prime}\cup\Gamma_{\psi_{2,1}}^{\prime}\cup\Gamma_{\psi_{2,2}}^{\prime}$.
@@ 3.3 Forward chaining
We now illustrate the graph construction and the corresponding FC procedure for the rule set $\Gamma_{\psi}=\{\psi_{1},\psi_{2}\}$ under the Level-2 interpretation given in Example 1. We first obtain all the literals as $\texttt{Lit}_{\mathcal{R}}=\{p_{0},p_{1},p_{2},p_{3},p_{4}\}$ with $\texttt{Lit}^{\uparrow}=\{p_{0},p_{2}\}$ and $\texttt{Lit}^{\downarrow}=\{p_{1}\}$, construct the literal nodes correspondingly, and construct the LHS node set as $\{p_{0}\&p_{2},p_{1}\&p_{2}\}$. Then, directed edges are added as follows: $p_{0}\rightarrow p_{0}\&p_{2}$, $p_{2}\rightarrow p_{0}\&p_{2}$, $p_{0}\&p_{2}\rightarrow p_{3}$, $p_{1}\rightarrow p_{1}\&p_{2}$, $p_{2}\rightarrow p_{1}\&p_{2}$, $p_{1}\&p_{2}\rightarrow p_{4}$, resulting an initial graph. Next, we check the logic consistency. For the LHS node $p_{0}\&p_{2}$, since $\{p_{0},p_{2}\}\subseteq\texttt{Lit}^{\uparrow}$, we update $\texttt{Lit}^{\uparrow}=\texttt{Lit}^{\uparrow}\cup p_{3}=\{p_{0},p_{2},p_{3}\}$ and $\texttt{Lit}^{\downarrow}=\texttt{Lit}^{\downarrow}\cup\neg p_{3}=\{p_{1},\neg p_{3}\}$, where $p_{3}={\sf True}$” is the new knowledge derived. For the LHS node $p_{1}\&p_{2}$, since $p_{1}\notin\texttt{Lit}^{\uparrow}$, there is no update on $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\uparrow}$, hence no new knowledge inferred.
@@ 3.4 Query generation
Clarification. The soundness of this work—the ability to detect inconsistencies with respect to domain-specific constraints—depends on the accurate and faithful encoding of domain knowledge by the expert. Thus, the validity of RvLLM hinges on the alignment between the specification given in ESL and the underlying domain knowledge, and any misrepresentations in this encoding may compromise the soundness of the method.
@@ 4 Experiments
To evaluate the effectiveness of RvLLM, we apply it to the runtime verification of LLMs across three representative tasks that require domain-specific knowledge: violation detection against Singapore Rapid Transit Systems Act [SGMRT], numerical comparison, and inequality solving. For a comprehensive evaluation, we include both SOTA and non-SOTA LLMs as our benchmark, including Qwen 2.5 (max, plus, turbo, 7B, 14B, 32B, 72B), GPT (4.1, 4.1 mini, 4.1 nano), Gemini 2.0 (Flash, Flash Lite), and DeepSeek-V3. All experiments are conducted on a machine with an Intel (R) Xeon(R) w7-2475X processor. A dedicated guideline for designing interpretation abstraction prompts for the perception agent is provided in the appendices, along with additional experimental details and descriptions of the datasets and specifications used throughout the experiments.
To ensure a comprehensive evaluation, we also compare RvLLM with other runtime approaches, including SelfCheckGPT [manakul2023selfcheckgpt_factual] and a prompt-based augmentation method that explicitly encodes domain-specific constraints within prompts. Due to the computational overhead and API costs associated with large-scale evaluations, we restrict our comparison to cost-efficient model variants. Additional experimental results and missing implementation details/justifications can be found in [zhang2025rvllm].
@@ 4.1 Case study 1: violation detection against Singapore Rapid Transit Systems Act
For this case study, we apply RvLLM under the Level-1 interpretation to evaluate the LLM’s responses in detecting violations within contextual scenarios against the Singapore Rapid Transit Systems Act. In this capacity, RvLLM can be regarded as a complementary mechanism to enhance the LLM’s capability in detecting the law violations in this study. We conduct experiments using contextual scenarios derived from [wang2024ali], consisting of 304 cases in total–281 labeled as involving violations (unsafe) and 23 as not (safe). To incorporate relevant domain knowledge, we carefully encode 31 of the 53 regulations from Singapore Rapid Transit System Act into ESL specifications. It took a research scientist three days to interpret the law and develop corresponding specifications–a one-time effort applicable for verifying any LLM application against these regulations. In this study, the same model serves both as the target LLM and as the perception agent.
@@
Figure 3 gives the comparison results of RvLLM against other baseline methods. We observe that SelfCheckGPT often yields the highest TPR gains, as it labels nearly all outputs as “hallucinations”, resulting in near-zero TNR. This outcome is expected since SelfCheckGPT relies solely on output stability without leveraging domain knowledge, making it more suitable for open-domain generation. The prompt augmentation method (also known as LLM-as-a-judge) achieves TPR performance comparable to RvLLM but accompanied by a non-negligible TNR decrease. In contrast, RvLLM achieves a better balance between TPR and TNR, yielding more reliable and interpretable results.
Table 1: Violation detection results of LLMs with and without RvLLM against the Singapore Rapid Transit Systems Act [SGMRT]. All percentage values are reported rounded to one decimal place.
@@
Qwen max [yang2024qwen2] 56.2% 91.3% 86.1% 91.3% +29.9% 0 3.00
Qwen plus [yang2024qwen2] 58.4% 1 81.9% 1 +23.5% 0 5.32
Qwen turbo [yang2024qwen2] 39.1% 1 74.7% 1 +35.6% 0 1.64
Qwen 2.5 (7B) [yang2024qwen2] 16.0% 1 61.2% 87.0% +45.2% -13.0% 2.28
Qwen 2.5 (14B) [yang2024qwen2] 38.8% 95.7% 80.4% 87.0% +41.6% -8.7% 3.11
Qwen 2.5 (32B) [yang2024qwen2] 56.6% 95.7% 87.5% 95.7% +31.0% 0 2.55
Qwen 2.5 (72B) [yang2024qwen2] 57.3% 1 80.4% 95.7% +23.1% -4.3% 2.57
GPT 4.1 [achiam2023gpt] 57.7% 95.7% 81.1% 91.3% +23.5% -4.3% 3.11
GPT 4.1 mini [achiam2023gpt] 39.1% 95.7% 65.1% 95.7% +26.0% 0 3.94
GPT 4.1 nano [achiam2023gpt] 11.7% 1 29.9% 1 +18.1% 0 1.82
Gemini 2.0 Flash [geminiteam2024gemini] 37.0% 1 87.2% 82.6% +50.2% -17.4% 2.69
Gemini 2.0 Flash Lite [geminiteam2024gemini] 54.8% 95.7% 79.0% 95.7% +24.2% 0 1.72
DeepSeek-V3 [liu2024deepseek] 78.3% 87.0% 94.0% 82.6% +15.7% -4.3% 15.84
(a) The increase of TPR across models.
(b) The decrease of TNR across models.
Figure 3: Performance comparison of different methods in the violation detection task, where QW, GF/GF-Lite, and DS-V3 denote Qwen, Gemini 2.0 Flash/Flash Lite, and DeepSeek-V3, respectively.
@@ 4.2 Case study 2: numerical comparison problems
Results. Table 3 gives the verification results by RvLLM on numerical comparison tasks across various LLMs. Columns Con. and Incon. report the number of cases where RvLLM detects no inconsistencies and where at least one inconsistency is detected, following the process in Figure 2. The values in parentheses in Column Con. represent cases where the target LLM produces an erroneous comparison result on the original comparison question, yet RvLLM fails to identify any inconsistencies–typically due to incomplete interpretation by the perception agents. In contrast, the values in parentheses in Column Incon. corresponds to cases where the target LLM returns a correct comparison result on the original question, but RvLLM indeed detects some inconsistencies. This usually arises from two main reasons: i) incomplete interpretation, and ii) the target LLM producing an incorrect diagnosis on newly-inferred knowledge. Overall, RvLLM effectively identifies erroneous diagnoses by the target LLM in most cases, with exceptions observed in only 3 cases for GPT 4.1 nano and mini. Among all tested models, Gemini 2.0 Flash Lite and DeepSeek-V3 show the weakest diagnostic performance, and Qwen 2.5 (7B) has the worst language understanding capabilities, failing all interpretation abstraction. Table 3 presents the performance comparison across methods. Consistent with the findings in case study 1, we find that RvLLM achieves the best balance, maintaining both high true positive and high true negative rates among all evaluated approaches.
Table 2: Verification results by RvLLM over various LLMs for numerical comparison tasks.
@@
Qwen max [yang2024qwen2] 93 7 0 5.03
Qwen plus [yang2024qwen2] 99 1 0 7.01
Qwen turbo [yang2024qwen2] 94 6(2) 0 2.93
Qwen 2.5 (7B) [yang2024qwen2] 0 0 100 N/A
Qwen 2.5 (14B) [yang2024qwen2] 90 10(7) 0 5.09
Qwen 2.5 (32B) [yang2024qwen2] 98 2 0 4.98
Qwen 2.5 (72B) [yang2024qwen2] 96 4 0 6.31
GPT 4.1 [achiam2023gpt] 98 2 0 4.84
GPT 4.1 mini [achiam2023gpt] 92(1) 8 0 5.37
GPT 4.1 nano [achiam2023gpt] 90(3) 10 0 3.73
Gemini 2.0 Flash [geminiteam2024gemini] 93 7 0 4.94
Gemini 2.0 Flash Lite [geminiteam2024gemini] 52 44 4 4.35
DeepSeek-V3 [liu2024deepseek] 69 31(5) 0 30.87
Table 3: Performance comparison of different methods for numerical comparison tasks.
Target LLMs Methods TPR TNR
SelfCheckGPT 0 96.7%
Qwen plus [yang2024qwen2] LLM-as-a-judge 100% 0
RvLLM 100% 100%
SelfCheckGPT 100% 0
GPT 4.1 mini [achiam2023gpt] LLM-as-a-judge 0 100%
RvLLM 88.9% 100%
SelfCheckGPT 87.5% 0
Gemini 2.0 Flash [geminiteam2024gemini] LLM-as-a-judge 0 100%
RvLLM 100% 100%
@@ 4.3 Case study 3: inequality solving problems
For this case study, we employ RvLLM under Level-1 interpretation to verify the step-by-step reasoning process of LLMs in inequality-solving tasks. We compile a dataset of 40 inequality questions sourced from A-Level H2 Mathematics examination papers [TestpaperJC] and carefully design three specifications to serve as domain knowledge that RvLLM uses for verification. These specifications consist of basic algebraic inequality properties taught at the junior college level, targeting three common LLM reasoning errors in inequality solving: incorrect factorization, flawed interval analysis, and omission of endpoint or critical point checks. Given the higher interpretation complexity in this case study and the strong language processing capabilities of DeepSeek-V3, we employ it as the perception agent for all LLMs evaluated and Qwen 2.5 (32B) as an alternative for comparison.
Results. Given the strong performance of RvLLM in detecting true positives, we focus the evaluation on the questions that the target LLMs answered incorrectly. The results are summarized in Table 4. Column 2 lists the total number of questions for which the target LLM produces incorrect solutions. Columns labeled DS-V3 and QW-32B (spanning Columns 3 to 11) show the number of cases where RvLLM successfully identifies inconsistencies based on the corresponding specifications using DeepSeek-V3 and Qwen 2.5 (32B) as the perception agent, respectively. Values in parentheses indicate false positives–cases where the detected inconsistency does not stem from an actual reasoning error but from inaccuracies of the perception agent during the interpretation abstraction process. Column GT gives the ground truth number of incorrect responses attributable to the rule specifications. It is important to note that we define only three specifications as domain knowledge for inequality-solving tasks, which is insufficient to cover all reasoning patterns required for such problems comprehensively. As a result, it is expected that RvLLM fails to detect a portion of the errors due to the lack of knowledge. The final two columns give the true positive rate of RvLLM with three specifications. We find that RvLLM achieves a TPR of up to 50% when using DeepSeek-V3 as the perception agent. This result underscores both RvLLM’s potential to detect reasoning errors—even with limited domain knowledge—and the critical role of a robust perception agent in boosting overall performance. Although TPR remains below 50% for most models, they rise markedly when restricted to ground truth cases (Column GT), suggesting substantial gains are possible with richer domain knowledge in inequality solving.
@@
Qwen max [yang2024qwen2] 12 1 1 1 1 0 1 2 2 3 33.3% 25%
Qwen plus [yang2024qwen2] 10 1 1 2 0 0 0 0 1 2 10% 20%
Qwen turbo [yang2024qwen2] 22 0 0 1 1 0 1 3 2 4 18.2% 9.1%
Qwen 2.5 (7B) [yang2024qwen2] 25 2 1 3 7 2 7 3 2 5 48% 20%
Qwen 2.5 (14B) [yang2024qwen2] 18 0 0 1 3 2(1) 3 3(1) 2 3 27.8% 16.7%
Qwen 2.5 (32B) [yang2024qwen2] 16 1 0 1 3 2 4 4 2 5 50% 25%
Qwen 2.5 (72B) [yang2024qwen2] 15 1 0 1 4 3 4 0 0 1 33.3% 20
GPT 4.1 [achiam2023gpt] 5 1 1(1) 1 0 0 0 0 0 0 20% 0
GPT 4.1 mini [achiam2023gpt] 2 0 0 0 0 0 0 1 0 1 50% 0
GPT 4.1 nano [achiam2023gpt] 10 0 0 0 1 0 1 1 0 1 20% 0
Gemini 2.0 Flash [geminiteam2024gemini] 5 0 0 0 1 0 1 0 0 0 20% 0
Gemini 2.0 Flash Lite [geminiteam2024gemini] 6 1 1 2 0 0 0 0 0 1 16.7% 16.7%
5 Related work
LLM testing. Research in this area aims to develop comprehensive benchmarks and systematic evaluation frameworks to assess LLM across general attributes such as robustness, fairness, and factual consistency [liu2023trustworthy, liang2022holistic, wang2023decodingtrust]. Robustness studies examine model resilience to adversarial input, paraphrasing, or noise [cai2024benchlmm, tu2023many, zhang2024multitrust], while fairness evaluations assess bias with respect to sensitive attributes like gender or race [li2024red, zhang2024multitrust]. Factual consistency detection, or hallucination detection, focuses on identifying instances in which LLMs produce plausible yet factually incorrect or unsupported claims [lin2022truthfulqa, li2023halueval, li2024drowzee]. Although large-scale benchmarks like HELM [liang2022holistic] and BIG-bench [srivastava2022beyond] offer task suites and metrics for performance evaluation, a critical challenge remains—the dynamic and open-ended nature of real-world deployments complicates the construction of exhaustive and representative test suites.
Runtime verification of LLMs. Runtime verification [bartocci2018introduction, falcone2013tutorial] is a dynamic analysis technique that ensures certified system behavior during execution by monitoring traces against formal specifications. Owing to its lightweight nature, runtime verification has inspired multiple efforts to verify LLM-generated outputs at runtime [cohen2023lm_factual, besta2024checkembed_factual, manakul2023selfcheckgpt_factual, chen2024complex_factual, laban2022summac_faithful, luo2023chatgpt_faithful, laban2023llms_faithful]. However, existing approaches center on open-domain settings, where specifications are often ambiguous and informal. For instance, SelfCheckGPT [manakul2023selfcheckgpt_factual] and CHECKEMBED [besta2024checkembed_factual] operationalize specifications in terms of output stability. More recently, [cheng2024formal] proposes a fairness monitoring approach leveraging precisely defined properties in formal specifications. These specifications, informed by linear temporal logic [pnueli1977temporal] and its bounded metric variant [alur1996benefits], signal a shift towards formal methods for more accurate and dependable runtime verification of LLM behaviors. However, these works primarily focus on general properties and are not well-suited to domain-specific constraints for specialized tasks.
Expert systems with NLP. Rule-based expert systems [daniel1997cadiag, denning1986towards, durkin1996expert] have effectively tackled domain-specific problems by coupling knowledge bases with inference engines to emulate expert reasoning for decades of years. To enhance usability, in the 70s, researchers integrated natural language processing (NLP), enabling interaction via simplified language, e.g., SHRDLU [winograd1971procedures] in virtual environments and MedLEE [krauthammer2001knowledge] in clinical text analysis. These systems relied on symbolic parsing and controlled vocabularies but suffered from fragile parsing, limited lexical coverage, and ambiguity, often requiring constrained input or extensive user guidance. Nevertheless, integrating NLP with rule-based reasoning provides a key proof-of-concept for interactive, user-friendly AI, shaping later knowledge-based systems and natural language interfaces. Recently, the rise of LLMs has reinvigorated this paradigm. In this work, we employ a perception LLM to translate the target LLM’s context and outputs into formal representations for backend reasoning. An improved version could explore other extensions of Horn clauses beyond those considered here to enhance expressiveness and completeness [babka2013complexity].